OpenStudy (anonymous):

What is the slope of line CD if C(5,-4) and D(5,2)? Is this a line or line segment?

OpenStudy (anonymous):

\(slope = \dfrac{rise}{run} = \dfrac{\text{change in y}}{\text{change in x}} =\dfrac{y_2 - y_1}{x_2-x_1}\)

OpenStudy (anonymous):

I got 0 is that correct?

OpenStudy (anonymous):

Yes, it is.

OpenStudy (anonymous):

And how do i know if its a line or a line segment?

OpenStudy (anonymous):

A line continues on forever in either direction. A line segment has endpoints.

OpenStudy (anonymous):

so it'll be a line segment bc it has an endpoint

OpenStudy (anonymous):

It seems that way. Usually in a word problem, you will be told if it is a line or a line segment. If there is a picture accompanying this question, a line will have arrows on either end.

OpenStudy (anonymous):

no theres no picture

OpenStudy (anonymous):

Is what you typed the exact wording of the question?

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

I'll type it again: What is the slope of CD if C(5,-4) and D(5,2)? Is this a line or a line segment?

OpenStudy (anonymous):

I would have to guess that it is a line segment. Maybe I am over-thinking this, but it just doesn't seem clear to me.